OpenStudy (anonymous):

5. Simplify the radical expression. Square root 20x^13y^5/5xy^7 A) square root 4x^12/y^2 B) 2x^6/y C) 2 square root x^12/y^2 D) 2x^6y

OpenStudy (anonymous):

again same concepts

OpenStudy (anonymous):

Start with \[\sqrt{20}\] To take this out of the radical, you need to find a squared term that's a factor of 20. What is that term?

OpenStudy (dannyo19):

None of the answers appear to be correct as is.. Are you sure about your syntax?

jhonyy9 (jhonyy9):

check them @DannyO19 newly because there is the right answer

jhonyy9 (jhonyy9):

commen this is easy

jhonyy9 (jhonyy9):

so sqrt20/5 = sqrt4 = 2 yes ?

OpenStudy (dannyo19):

yes

jhonyy9 (jhonyy9):

sqrtx^13/x = sqrtx^12 = x^6 yes ?

OpenStudy (dannyo19):

Yeah, sure

jhonyy9 (jhonyy9):

and sqrty^5/y^7 = sqrt(1/y^2) = 1/y so tha right answer will be B. ok. ?

jhonyy9 (jhonyy9):

do you understand it @Answers101 ?

OpenStudy (dannyo19):

No, that's not right

jhonyy9 (jhonyy9):

yes but in denominator are x and y^7 yes ?

jhonyy9 (jhonyy9):

do you can rewrite this exercise without denominator for you can check it sure ?

jhonyy9 (jhonyy9):

with exponents negativ - this is easy - try it

OpenStudy (dannyo19):

The problem is that once we factor out the 2 we get the Sqrt of ( (x^13y^5)/(xy^5) ) an expression that can't be further simplified because the numerator and denominator aren't like terms.

jhonyy9 (jhonyy9):

square root 20x^13y^5/5xy^7 = sqrt(20x^13y^5*5^(-1)*x^(-1)*y^(-7) = =sqrt(4*5*5^(-1)*x^(13-1)*y^(5-7) ) = 2sqrt(x^12*y^(-2) ) = (2x^6)/y so this mean that B is right answer sure

jhonyy9 (jhonyy9):

@Answers101 do you understand it why B. is the right answer ?

jhonyy9 (jhonyy9):

@mathmale do you understand what have said @Danny ?

jhonyy9 (jhonyy9):

@ganeshie8 please your opinion about this ?

OpenStudy (mathmale):

5. Simplify the radical expression. Square root 20x^13y^5/5xy^7 do you mean the following? See below \[\frac{ 20x ^{13}*y^5 }{ 5xy^7 }??\] A) square root 4x^12/y^2 B) 2x^6/y C) 2 square root x^12/y^2 D) 2x^6y

OpenStudy (mathmale):

Please either verify this, or correct it.

jhonyy9 (jhonyy9):

@mathmale who have said you this above ?

OpenStudy (mathmale):

That comment was directed at Answers101.

jhonyy9 (jhonyy9):

ok. thanks